OpenStudy (anonymous):

Complete the square below to solve the equation below. Check all that apply. x^2+8x-5=15 A.X=-10 C.X=-3 D.X=-2

jimthompson5910 (jim_thompson5910):

x^2+8x-5=15 x^2+8x-5+5=15+5 x^2+8x = 20 what must you add to both sides to complete the square?

OpenStudy (anonymous):

20?

jimthompson5910 (jim_thompson5910):

you were doing it with the last 5 or so problems

OpenStudy (anonymous):

ohh my bad 8

jimthompson5910 (jim_thompson5910):

you add 8 to both sides?

jimthompson5910 (jim_thompson5910):

look back at your previous posts

OpenStudy (anonymous):

half of 8=4 so +4

jimthompson5910 (jim_thompson5910):

now square that

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

so you add that to both sides

OpenStudy (anonymous):

x=-10 right

jimthompson5910 (jim_thompson5910):

good, what else

OpenStudy (anonymous):

x^2+2x-9=15